Run Planning Meeting Notes, April 3 - April 9, 2025
From GlueXWiki
The zoom link for the Hall D 0845 meeting is the first in this list
Hall D counting room whiteboard
Back to Run Coordination Meetings:Spring2025#Runplan
Contents
Updates from Wednesday April 8 morning
- Run 130889 - ADC & TDC times for TOF are shifted. Peak moved 130 ish to 25. TDC at 90 instead of 190. Something shifted, Beni does not know what - most likely the trigger. Since the TOF timing is bad, most likely the other detectors are bad too.
Naomi will ask Sasha if he knows the explanation. He will be updating the trigger config files today.
Wednesday April 8
Last 24h
- Maintenance day - down for day shift and first half of swing
- Sasha provided an updated daq config including BCAL. Untested, so have earlier config as a fallback. (but it did not include BCAL)
- Many issues were fixed: BOR, run numbers in ELog entries, halld_recon was rebuilt including ET and file-copying to tape resumed.
- Again, the overnight runs used only ECAL triggers.
- Shift summaries: day swing owl
Reports
- TPOL missing channels restored? (Beni) Fixed! but not in the CH monitoring plots (do we need an update?)
- Also unswapped some TOF channels, also have to wait and see (Beni)
- ECAL work (Sasha)
- First look at FCAL hardware gains (Malte)
- New harp scan extrapolation (Phoebe)
- Active Collimator remained unlocked after harp scan, reinstated at 4am (Phoebe)
Problems
- ECAL crate missing from readout since the start looks ok from Sasha's separate daq config
- The DAQ behaved better overnight - no issues with livetime dropping after a few minutes, but ECAL5 hung when the run was ended.
- Request for automatic HDRUN entries
- PS DAC values may be off (Beni)
- Request to add printer to gluon02 -> Sergey investigated and asked Paul Letta to fix
- Need to update BTA guidelines as the shifters can now modify ACC and OFF columns (new!)
Plans
- 5th pass RF separator commissioning - down from approx 0830 until 1500/1600 - we can access the hall if we need to.
- (They need to phase each of the 4 cavities to get them in sync to do the separation. haven't done this exact process before. In the future it should take about half a shift.)
- On Friday it will take a few h to configure the 5th pass vertical separator for Hall B. Expect a few h downtime.
- DAQ configuration requests
- Add ECAL5 to the readout
- Include BCAL and maybe FCAL in the trigger
Tuesday April 8
Last 24h
- Sasha: low energy TAGH counters and ECAL are on, we should stay at 100 nA with 4.5E4 RL radiator for now.
- Beni's text reported beam for 11 of the last 24h.
- TAGM work completed (Richard). This means beamline setup is complete. EDIT: not quite, Richard still has to calculate the settings offline and load them.
- BTA (BTM) issue fixed (Hovanes)
- RCDB's temporary website whitelisted :-) (Edmundo)
- Raw mode run with BCAL trigger 130832 (Sergey)
- Production mode runs with BCAL trigger 130833-7
- Sasha made an access to replace a HV module for ECAL
- Sasha updated the trigger configuration to include ECAL. Latest DAQ config described here
- Lubomir took drift chamber scan data.
- Continued taking data through the night, but had a variety of problems with the DAQ and other software.
- Shift summaries: day swing owl
Reports
- Radiation levels (Alexandre)
- BPM calibration (MCC)
- Timing updates (Sean)
- ECAL scalers ECAL trigger commissioning run started (Sasha)
Problems
- The DAQ is not ok. Problems through the night
- No rootspy! halld_recon issue?
- DAQ comments do not reach RCDB
- EVIO data not copied to cache since 11pm yesterday (Beni)
- BOR missing (Sean)
- Run number not updating in logbook screenshot tool (Justin)
- and https://logbooks.jlab.org/entry/4350948 TOF issues (Beni)
- Hot channels in PS fine counters left arm fADC channels (Justin)
- Missing TPOL hits: sectors 23-25 (Justin) -> Mike believes this is mis-cabling that we should fix during access
Plans
- MCC news: Should have beam back to us mid-swing. A getting beam end of today. B -> 5th pass on Friday? if all is well.
- Wednesday: Commissioning improved RF separator, starting ~ 0830 and likely to take the full day shift. Also C -> 5th pass.
- Maintenance day - controlled access during day shift. Remember to close the doors properly!
- Day shift worker on standby until 4pm. Shift leader will be on duty.
- Hall walk-through for training at 1pm - to attend, contact Mark
- Resume ECAL commissioning when beam returns
- TOF HV scans later on (Beni)
Monday April 7
Updates from the weekend
Friday
- Shifts cancelled. Saturday owl, duck, day on standby
- 11am A leak of environmentally hazardous water from the accelerator was found, and fixed.
- 2pm Magnet issue was found, and fixed.
Saturday
- 10am harp scan looked great!
- Shifts resumed at noon.
- Vacuum pump stopped in the tagger hall. Scot mended it (I think he replaced a control board).
- 3pm beam tuning started, continued intermittently through swing and into owl shifts (2 issues w RF cavities caused delays)
- Shift summaries: day swing
Sunday
- During owl shift, the ion chamber calibrations and some radiation levels checks were completed.
- Early day shift, Hovanes arrived to commission the new profiler. This went very well until he left. The beam was locked on the profiler and all looked good; MCC took the beam away to set up hall B. When beam was restored some hours later, the profiler readings were extremely erratic, so we switched the lock back to 5C11B. Turns out that this was due to not quite enough HV on the profiler.
- Richard set up the A.C. We can now lock on 5C11B and AC.
- Radiation level checks were completed.
- Sasha switched off the low E TAGH counters and provided some initial DAQ config files.
- Hovanes completed the profiler scan.
- Richard started the TAGM setup. Should be finished by 9 ish. Monday
- Shift summaries: owl day swing Monday owl
Reports
- harp scan extrapolation plots, Sunday morning
- 130769 monitoring plots can be seen in plot browser - first run with BCAL trigger, took data throughout the radiation studies, so it contained a variety of beam intensities. A few hot channels in the calorimeters, ST, some plots missing.
- Profiler with raised HVs
Problems
- BTA epics script not working - Hovanes requested the old one back. Naomi has to correct some entries. MCC person provided Hovanes with some info to fix this.
- The DAQ config is NOT ok - there were problems with rocs at the end of the run. Richard's logentry
- RCDB website is not accessible from the gluon browsers, blocked by firewall. Have to use laptop.
- Rootspy issues - TRD histo was not working properly, eLog entries need to specify the run number.
- Shift primer is out of date - JTAB menu has gone.
- Everyone needs reminders to read the hall safety docs and sign the sheet.
Plans
- MCC: Tues maintenance. Starting 7? Will close ~ 2:30pm. Swing: set up A for beam, maybe C to 5th pass (might postpone to Wed day). B to 5th pass maybe Fri or Mon. C needs arc energy measurement, should not interrupt our beam.
- Monday: TAGM setup is continuing, need to do ECAL/FCAL commissioning, detector checkout, detector HV scans
- Tuesday: maintenance day. We asked LBL parasitic experimenters to be here at 0730 to be ready to make their 30 min access before HD people need it. List for opportunistic access is on the whiteboard. Mark will lead a walkthrough at 1pm.
Friday April 4
Past 24h
- Harp problem caused no beam - fix was attempted but found to have failed. MCC eventually abandoned attempts to steer beam without it, need to consult experts.
- Hall D had to be swept because the outer door was not closed properly. PSS box "Lockdown/PSS safe". Maybe it bounced back after slamming.
- Persistent disconnect alarms from switched off crates are a nuisance. Please inform shift-takers.
Thursday April 3
Past 24h: no beam.